# A Lot of Talk and a Badge: An Exploratory Analysis of Personal Achievements in GitHub

### 30 Dec 2023 v2

@@ 4.2 Methods
To answer RQ (diffusion of personal badges), we compute the number and percentages of badges unlocked by developers in our experimental sample for each month, from June to December 2022, together with the number of developers who opted out of visualizing their badges on the profile page or made their profile page private. In addition, we plot the monthly progression of the number of unlocked badges.
Regarding the RQ (reliability of badges as signals), we first hypothesize and then search for correlations between the presence of personal badges and differences in the qualities they might signal. We collect several measures, such as the number of followers, commits, and open and closed issues and PRs. After verifying the presence of non-normal distributions with the Shapiro-Wilk test, we use the non-parametric Wilcoxon-Mann-Whitney (WMW) test to compare distributions and Cliff’s to assess the effect size.
@@ 5.1 RQ – Diffusion of Personal Badges
Table 1 shows the distribution of unlocked badges in our data set as of December 2022. The statistics in the table reveal that the most popular badge is Arctic Code Vault Contributor (4591, 27.02%). This result is ‘inflated’ because this one is among the badges that already existed before the launch of the personal badges in June 2022; besides, this badge is not earnable anymore as it was awarded to anyone contributing code to GitHub in 2020. The other previously existing badges are Mars 2020 Contributor (282, 1.66%), and Public Sponsor (118, 0.69%), unlocked respectively by developers who contributed code to the repositories used in the Mars 2020 helicopter mission and by those who have provided sponsorships to projects via the GitHub Sponsors program. Because these three badges predate the introduction of personal achievement badges in June 2022, they were excluded from subsequent analyses.
Regarding the other recently introduced types of badges, the most common (aggregated by tier) is Pull Shark (4,762, 28.03%), unlocked by merging a certain number of pull requests. The second most common type of badge in our dataset is Starstruck (2,385, 14.03%), which is awarded by owning repositories that receive more and more stars from other users. Pair Extraordinaire (1,824, 10.74%) and Galaxy Brain (263, 1.55%) are unlocked respectively by coauthoring merged pull requests and by answering questions in project discussions. Finally, YOLO (1,324, 7.79%) is the personal badge awarded to those who have merged a pull request without performing a code review. Because the action that earns developers this badge is a development practice that should not be encouraged, it was also excluded from subsequent analyses. We also discuss the bad coding practice promoted by the YOLO badge during the analysis of the survey responses in Section 5.4.1.
As of December 2022, most of the 6,022 developers in our dataset had at least one personal badge displayed on their profile page (4,977, 82.65%), only 1 (0.02%) had no personal badge, and 1,044 (17.34%) had opted out of displaying badges or having a public profile. Figure 2 plots over a six-month time span the distributions of the number of developers with badges (with pre-existing badges excluded) versus those who either opted out of displaying badges on their profile page or without a public profile altogether. We observe an initial increase in the number of users who display badges from June (4,825, 80.12%) to July (5,163, 85.74%); afterward, we notice a decreasing trend (4,452, 73.93% in November) with more and more developers choosing not to display badges, either opting out or because they chose not to have a public profile page (1,412, 23.48% in November).
@@ 5.2 RQ – Reliability of Personal Badges as Signals
To gauge how GitHub developers perceive the signals sent by personal badges, we defined the following question in the survey (further described in Section 5.4): “Q. For each badge, please report the intended signal that you think it sends to others and, if different, the actual signal it conveys to people visualizing it.” As detailed in Section 5.4.1, 54 GitHub developers participated in the survey. However, for question Q we received 31 answers. The somewhat low number of answers suggests that, in general, the GitHub community may not have a clear understanding of what signals, if any, personal badges send to others; as P wrote: “they’re supposed to show me some kind of evidence, but I don’t know.” Consistently, in many cases the respondents replied that they had no clue about the meaning of some badges; however, when they did, their answers showed some consistency in describing the perceived signals. After analyzing and grouping the mentions of similar signals in the answers, which we considered only if they appeared at least three times, we hypothesized associations between the perceived signal sent by each personal badge and some code development metrics available through the GitHub API, as described next.
@@
The Pair Extraordinaire badge is awarded to developers who coauthor merged pull requests. According to the survey responses (), owning the badge might signal an increased desire to collaborate (: “[the badge] suggests that they like collaborating with other developers to create some pieces of code”). Accordingly, we tested its correlation with the number of authored commits, as well as the number of PRs and Issues opened, as pull-based development has been popularized by GitHub to collaborate and integrate developers’ contributions to project repositories (Gousios et al., 2015, 2016); also, we aggregated the number of PRs and Issues opened, due to the interlinked nature in GitHub and because opening an Issue (e.g., for reporting bugs or requesting features) is a valuable form of non-coding collaboration (Izquierdo and Cabot, 2022). The results of the WMW tests show the existence of a statistically significant difference between developers with and without the badge and the number of opened Issues and PRs (Figure 2(a), ) with a negligible effect size (). Instead, there is no statistically significant difference in the case of the number of authored commits between developers with and without the Pair Extraordinaire badge (Figure 2(b)).
@@
Regarding the Quickdraw badge, gained by developers who close Issues and PRs within 5 minutes, according to the survey response () its presence could be a signal of timeliness (: “it may indicate that you’re fast, committed;” : “The word [Quickdraw] says is it… they’re quick;” : “I guess it signals you’re on time, which means dedication too”). Consequently, we tested its association with an overall shorter time taken to close Issues and PRs compared to those who do not own the badge. The result (Figure 2(c)) shows a statistically significant difference in the WMW test but a negligible effect size (, ).
@@
The most substantial finding is observed for the Starstruck badge, which, according to the respondents () is perceived as a signal of popularity since it is unlocked when repositories are starred (: “It tells others that you are a coding rockstar;” : “[It] means one is popular, I guess;”); therefore, we tested the correlation of its presence on developers’ profile pages with the number of their followers, a commonly used proxy measure of developer popularity in GitHub (Blincoe et al., 2016); the results in Figure 2(d) show that there is a statistically significant difference between the distributions of developers with and without the badge in terms of the number of followers () and that this difference has a large effect size ().
@@
The Galaxy Brain badge is unlocked when the answers posted to project discussions are accepted. According to respondents (), its presence could signal an increased willingness to help projects grow (: “Maybe support? That you want to help a GitHub project and its community beyond ‘just’ coding?”). Discussions (Hata et al., 2022) is a new feature for asking questions or discussing topics in GitHub projects, outside of specific Issues or PRs, in a fashion similar to Question-Answering sites such as Stack Overflow. The analogy with Stack Overflow suggests that developers may be motivated to answer posts in GitHub for similar reasons, most notably the desire to help fellow programmers (Calefato et al., 2018), contribute to the community (Lu et al., 2022), and build a reputation (Bosu et al., 2013). As such, we tested the correlation between its presence and the number of opened Issues and PRs, which again are the main features of GitHub used by developers to contribute to projects (Gousios et al., 2016). The results in Figure 2(e) show that there is a statistically significant difference () and a small effect size ( between developers with and without the badge in terms of opened Issues and PRs.
@@
Finally, the Pull Shark badge is obtained when PRs are merged into a project repository. Based on the survey responses (), its presence could represent a signal of willingness to contribute code to projects (: “As a maintainer, [Pull Shark] gives me the general idea of someone who wants to write code, for their own or others’ repos;” : “[It] signals that you’re an active code contributor”). Consequently, we tested its correlation with the number of both authored and committed commits. In the case of authored commits, the results of the WMW tests show a statistically significant difference, but a negligible effect size between developers with and without the badge (Figure 2(f), , ). Regarding committed commits, we observe a lack of statistical significance (see Figure 2(g)).
@@
Correlation analysis by tier. In general, the correlation analysis revealed limited significant results (see Table 2, column All). Most differences, in fact, are significant but with negligible or small effect sizes, with the exception of the Starstruck badge, which exhibits a strong association with the number of followers, as a signal of popularity. One reason for the negligible effect sizes may be due to having conducted the correlation analysis by aggregating all badge tiers (see Table 1), thus distinguishing only between developers with and without badges. Therefore, it might be possible that the tier-2 (bronze), tier-3 (silver), and tier-4 (gold) badges, which are harder to achieve, might send more reliable signals than the basic tier-1 badges of the same type. Consequently, in Table 2, we replicate the correlation analyses by tier. However, we observe only a few variations. We found that the difference in the number of followers between developers without the Starstruck badge and those who unlocked the basic tier is significant but with a negligible effect size (); on the contrary, the effect sizes remain large for the other tiers. Furthermore, for the Galaxy Brain badge, we observe a significant difference with a medium effect size () only for the silver tier.
@@ 5.3 RQ – Effects of Personal Badges on Developers’ Activities
This section presents the results of the difference-in-differences (DiD) regression analysis. The fundamental assumption for applying a DiD regression is that in the absence of treatment (i.e., before the introduction of personal badges) the dependent variable trend would be the same in both the treatment group (i.e., the developers who will eventually unlock a given type of badge) and the control group (i.e., the developers without the badges). In Figure 3(a), through the boxplots we visually confirm the parallel (or common) trends assumption for the Starstruck badge; Figure 3(b) shows instead a violation of the premise for the Quickdraw badge, which is therefore excluded from this analysis. The other badge excluded from this analysis for the same reason is Pull Shark, in terms of both the number of authored and committed commits. Refer to C for the other figures.
@@
Table 3 reports the results of the DiD regressions. They show that the interaction term is significant only for the Starstruck badge, on which we focus on the presentation of findings. The DiD coefficient is significant and different from zero (0.169); this means that the introduction of the new feature of the Starstruck badge, which counts the number of starred repositories of a developer, caused an increase in the average number of followers; specifically, because the regression is in logs, the average count of followers in the post-treatment time window is 18.4% (i.e., ) higher than it would have been without the introduction of the Starstruck badge in GitHub (Cameron and Trivedi, 2005). In addition, we observe that the coefficient (treatment group) is significant and different from zero, which means that in the pre-treatment time window the developers in the treatment and control groups had a different number of followers; specifically, developers with the badge had on average 299.9% more followers than those without. Finally, the coefficient is also significant and different from zero (-0.1), meaning that the average number of followers in the control group from the pre-treatment to the post-treatment time window decreased by 9.5%.
@@ 5.4.1 Survey analysis
We received 57 responses. Survey analysis was carried out by the first two authors. They analyzed the survey results reading and annotating the responses in a spreadsheet to extract interesting excerpts from open-ended questions and highlight commonalities. After that, the entire team discussed and consolidated the extracted excerpts.
@@
Regarding the demographics (Q-Q), participants in the survey reported having on average 9 years of experience (min. 1, max. 31). The participants also maintain an average of 6 open source software repositories (min. 0, max. 18) and have contributed to 22 repositories (min. 1, max. 100). Finally, they reported having an average of 28 followers (min. 0, max. 100). Three respondents were disqualified because they reported (Q) that they had opted out of displaying badges (1) or did not have a public profile page (2). They motivated their choices (Q) by saying that they “simply don’t like the idea of badges” (P) and that GitHub “may be a social coding site but it’s not (or shouldn’t be or become) like social media” (P). The other users who reported keeping the badge feature enabled (Q, ), cited various motivations for doing so. The primary reason was curiosity (19 respondents), followed by having observed badges on other developers’ profiles (18 respondents), the aesthetic appeal (12), as a suggestion by fellow GitHub users (3), and in recognition of the feature’s utility in conveying valuable information (2). In the following, we analyze the remaining valid responses.
@@
Question Q (, see Figure 5) aimed to assess whether participants consider the presence of personal badges an indicator of their own coding skills in general. Most of the respondents either strongly disagreed ( or disagreed (. Then, questions Q-Q () asked participants to report on the effects that badges may have on collaboration with other developers. The results are also reported in Figure 5. Although the distribution of responses is varied, overall the figure shows that about half of the respondents feel that badges can be neither an indicator of others’ ability nor are they influenced by their presence during collaboration with others. Question Q asked the respondents to justify their previous responses. They were critical of the current implementation of badges. Participant P found that “these badges can be achieved too easily, not like in Stack Overflow,” and P that there are too few of them (“we all have almost the same badges”); others added that they fail to adequately capture developers’ experience – e.g., “I have 30+ years of development experience and almost none is visible through [badges]” (P), “I have lots of experience and almost no badges” (P). Furthermore, P answered with a provocative question: “Are developers with these badges better developers? Vice versa and more critical: is a developer bad just because they don’t have some of these badges?”
@@
Next, we examine how respondents feel about specific badges. Question Q () asked participants to indicate the most relevant badges. The answers are clustered around two main groups. The first group contains answers (15) from users who feel that none of these badges can tell us anything useful about others. The other group of remaining answers (26) instead believes that the badges give information about who unlocks them. Specifically, 6 respondents mentioned the badge Galaxy Brain (awarded to those who have accepted answers in project discussions) because it reflects “how serious [one is] in the community” (P). A group of 11 participants mentioned the Starstruck badge (unlocked by receiving project stars) because it gives tangible “evidence of acknowledgment and interest by the community” (P). The remaining respondents (9) highlighted the importance of Pair Extraordinaire (awarded to those who coauthor merged pull requests) because the badge may reflect “a positive attitude [toward] collaboration with others” (P).
@@
Finally, question Q () asked participants to indicate any missing badge, in addition to those available. The suggestions converged on the following ideas for new badges that would display: i) the years of work in a project, ii) how many repositories you are a maintainer of, iii) how many repositories you have contributed to, iv) more badges in general, and v) more badges as long as they are challenging to unlock.

### 02 Feb 2024 v3

@@ 4.2 Methods
To answer RQ (diffusion of personal badges), we compute the number and percentages of badges unlocked by developers in our experimental sample for each month, from June to December 2022, together with the number of developers who opted out of visualizing their badges on the profile page or made their profile page private. Also, we plot the monthly progression of the number of unlocked badges.
Regarding RQ (reliability of badges as signals), we first hypothesize and then search for correlations between the presence of personal badges and differences in the qualities they might signal. We collect several measures, such as the number of followers, commits, and open and closed issues and PRs. After verifying the presence of non-normal distributions with the Shapiro-Wilk test, we use the non-parametric Wilcoxon-Mann-Whitney (WMW) test to compare distributions and Cliff’s to assess the effect size.
@@ 5.1 RQ – Diffusion of Personal Badges
Table 1 shows the distribution of unlocked badges in our data set as of December 2022. The statistics in the table reveal that the most popular badge is Arctic Code Vault Contributor (4591, 27.02%). This result is ‘inflated’ because it is among the badges that already existed before the launch of the personal badges in June 2022; besides, this badge is not earnable anymore as it was awarded to anyone contributing code to GitHub in 2020. The other previously existing badges are Mars 2020 Contributor (282, 1.66%), and Public Sponsor (118, 0.69%), unlocked, respectively, by developers who contributed code to the repositories used in the Mars 2020 helicopter mission and by those who have provided sponsorships to projects via the GitHub Sponsors program. Because these three badges predate the introduction of personal achievement badges, they were excluded from subsequent analyses.
Regarding the other recently introduced types of badges, the most common (aggregated by tier) is Pull Shark (4,762, 28.03%), unlocked by merging a certain number of pull requests. The second most common type of badge in our dataset is Starstruck (2,385, 14.03%), which is awarded by owning repositories that receive more and more stars from other users. Pair Extraordinaire (1,824, 10.74%) and Galaxy Brain (263, 1.55%) are unlocked, respectively, by coauthoring merged pull requests and by answering questions in project discussions. Finally, YOLO (1,324, 7.79%) is the personal badge awarded to those who have merged a pull request without performing a code review. Because the action that earns developers this badge is a development practice that should not be encouraged, it was also excluded from subsequent analyses. We also discuss the bad coding practice promoted by the YOLO badge during the analysis of survey responses in Section 5.4.1.
As of December 2022, most of the 6,022 developers in our dataset had at least one personal badge displayed on their profile page (4,977, 82.65%), only 1 (0.02%) had no personal badge, and 1,044 (17.34%) had opted out of displaying badges or having a public profile. Figure 2 plots over six months the distributions of the number of developers with badges (with preexisting badges excluded) versus those who either opted out of displaying badges on their profile page or without a public profile altogether. We observe an initial increase in the number of users who display badges from June (4,825, 80.12%) to July (5,163, 85.74%); afterward, we notice a decreasing trend (4,452, 73.93% in November) with more and more developers choosing not to display badges, either opting out or because they chose not to have a public profile page (1,412, 23.48%).
@@ 5.2 RQ – Reliability of Personal Badges as Signals
To gauge how GitHub developers perceive the signals sent by personal badges, we defined the following question in the survey (further described in Section 5.4): “Q. For each badge, please report the intended signal that you think it sends to others and, if different, the actual signal it conveys to people visualizing it.” As detailed in Section 5.4.1, 54 GitHub developers participated in the survey. However, for question Q, we received 31 responses. The somewhat low number of answers suggests that, in general, the GitHub community may not have a clear understanding of what signals, if any, personal badges send to others; as P wrote: “they’re supposed to show me some kind of evidence, but I don’t know.” Consistently, in many cases, the respondents replied that they had no clue about the meaning of some badges; however, when they did, their answers showed some consistency in describing the perceived signals. After analyzing and grouping the mentions of similar signals in the answers, which we considered only if they appeared at least three times, we hypothesized associations between the perceived signal sent by each personal badge and some code development metrics available through the GitHub API, as described below.
@@
The Pair Extraordinaire badge is awarded to developers who coauthor merged pull requests. According to the survey responses (), owning the badge might signal an increased desire to collaborate (: “[the badge] suggests that they like collaborating with other developers to create some pieces of code”). Consequently, we tested its correlation with the number of authored commits, as well as the number of PRs and Issues opened, as pull-based development has been popularized by GitHub to collaborate and integrate developers’ contributions to project repositories (Gousios et al., 2015, 2016); also, we aggregated the number of PRs and Issues opened, due to their interlinked nature in GitHub and because opening an Issue (e.g., for reporting bugs or requesting features) is a valuable form of non-coding collaboration (Izquierdo and Cabot, 2022). The results of the WMW tests show the existence of a statistically significant difference between developers with and without the badge and the number of Issues and PRs opened (Figure 2(a), ), but with a negligible effect size (). Instead, there is no statistically significant difference in the case of the number of authored commits between developers with and without the Pair Extraordinaire badge (Figure 2(b)).
@@
Regarding the Quickdraw badge, gained by developers who close Issues and PRs within 5 minutes, according to the survey response () its presence could be a signal of timeliness (: “it may indicate that you’re fast, committed;” : “The word [Quickdraw] says is it… they’re quick;” : “I guess it signals you’re on time, which means dedication too”). Consequently, we tested its association with an overall shorter time taken to close Issues and PRs compared to those who do not own the badge. The result (Figure 2(c)) shows a statistically significant difference in the WMW test, but a negligible effect size (, ).
@@
The most substantial finding is observed for the Starstruck badge, which, according to the respondents () is perceived as a signal of popularity since it is unlocked when repositories are starred (: “It tells others that you are a coding rockstar;” : “[It] means that you’re popular, I guess;”); therefore, we tested the correlation of its presence on developers’ profile pages with the number of their followers, a commonly used proxy measure of developer popularity in GitHub (Blincoe et al., 2016); the results in Figure 2(d) show that there is a statistically significant difference between the distributions of developers with and without the badge in terms of the number of followers () and that this difference has a large effect size ().
@@
The Galaxy Brain badge is unlocked when answers posted to project discussions are accepted. According to respondents (), its presence could signal a greater willingness to help projects grow (: “Maybe support? That you want to help a GitHub project and its community beyond ‘just’ coding?”). Discussions (Hata et al., 2022) is a new feature for asking questions or discussing topics in GitHub projects, outside of specific Issues or PRs, in a fashion similar to Question-Answering sites such as SO. The analogy with SO suggests that developers may be motivated to respond to GitHub posts for similar reasons, most notably the desire to help fellow programmers (Calefato et al., 2018), contribute to the community (Lu et al., 2022), and build a reputation (Bosu et al., 2013). As such, we tested the correlation between its presence and the number of Issues and PRs opened, which again are the main features of GitHub used by developers to contribute to projects (Gousios et al., 2016). The results in Figure 2(e) show that there is a statistically significant difference () and a small effect size ( between developers with and without the badge in terms of opened Issues and PRs.
@@
Finally, the Pull Shark badge is obtained when PRs are merged into a project repository. Based on survey responses (), its presence could represent a signal of willingness to contribute code to projects (: “As a maintainer, [Pull Shark] gives me the general idea of someone who wants to write code, for their own or others’ repos;” : “[It] signals that you’re an active code contributor”). Consequently, we tested its correlation with the number of both authored and committed commits. In the case of authored commits, the results of the WMW tests show a statistically significant difference, but a negligible effect size between developers with and without the badge (Figure 2(f), , ). Regarding committed commits, we observe a lack of statistical significance (see Figure 2(g)).
@@
Correlation analysis by tier. In general, the correlation analysis revealed limited significant results (Table 2, column All). Most differences, in fact, are significant but with negligible or small effect sizes, with the exception of the Starstruck badge, which exhibits a strong association with the number of followers, as a signal of popularity. One reason for the negligible effect sizes may be that the correlation analysis was performed by aggregating all badge tiers (Table 1), thus distinguishing only between developers with and without badges. Therefore, it might be possible that the tier-2 (bronze), tier-3 (silver), and tier-4 (gold) badges, which are harder to achieve, might send more reliable signals than the basic tier-1 badges of the same type. Consequently, in Table 2 we report the results of the correlation analysis by tier. However, we observe only a few variations. The difference in the number of followers between developers without the Starstruck badge and those who unlocked the basic tier is significant, but with a negligible effect size (); on the contrary, the effect sizes remain large for the other tiers. Furthermore, for the Galaxy Brain badge, we observe a significant difference with a medium effect size () only for the silver tier.
@@ 5.3 RQ – Effects of Personal Badges on Developers’ Activities
This section presents the results of the difference-in-differences (DiD) regression analysis. The fundamental assumption for applying a DiD regression is that in the absence of treatment (i.e., before the introduction of personal badges) the dependent variable trend would be the same in both the treatment group (i.e., the developers who will eventually unlock a given type of badge) and the control group (i.e., the developers without the badges). In Figure 3(a), through the boxplots we visually confirm the parallel (or common) trends assumption for the Starstruck badge; Figure 3(b) shows instead a violation of the premise for the Quickdraw badge, which is therefore excluded from this analysis. The other badge excluded from this analysis for the same reason is Pull Shark, in terms of both the number of authored and committed commits. Refer to the supplementary material in the replication package for the other figures.
@@
Table 3 reports the results of the DiD regressions. They show that the interaction term is significant only for the Starstruck badge, on which we focus on the presentation of findings. The DiD coefficient is significant and different from zero (0.169); this means that the introduction of the new feature of the Starstruck badge, which counts the number of starred repositories of a developer, caused an increase in the average number of followers; specifically, because the regression is in logs, the average count of followers in the post-treatment time window is 18.4% (i.e., ) higher than it would have been without the introduction of the Starstruck badge in GitHub (Cameron and Trivedi, 2005). In addition, we observe that the coefficient (treatment group) is significant and different from zero, which means that in the pre-treatment time window the developers in the treatment and control groups had a different number of followers; specifically, the developers with the badge had on average 299.9% more followers than those without it. Finally, the coefficient is also significant and different from zero (-0.1), meaning that the average number of followers in the control group from the pre-treatment to the post-treatment time window decreased by 9.5%.
@@ 5.4.1 Survey analysis
We received 57 responses. The survey analysis was carried out by the first two authors. They analyzed the survey results reading and annotating the responses in a spreadsheet to identify interesting excerpts from open questions and highlight commonalities. After that, the entire team discussed and consolidated the extracted excerpts.
@@
Regarding demographics (Q-Q), participants in the survey reported having on average 9 years of experience (min. 1, max. 31). The participants also maintain an average of 6 open source software repositories (min. 0, max. 18) and have contributed to 22 repositories (min. 1, max. 100). Finally, they reported having an average of 28 followers (min. 0, max. 100). Three respondents were disqualified because they reported (Q) that they had opted out of displaying badges (1) or did not have a public profile page (2). They motivated their choices (Q) by saying that they “simply don’t like the idea of badges” (P) and that GitHub “may be a social coding site but it’s not (or shouldn’t be or become) like social media” (P). The other users who reported keeping the badge feature enabled (Q, ), cited various motivations for doing so. The primary reason was curiosity (19 respondents), followed by having observed badges on other developers’ profiles (18 respondents), the aesthetic appeal (12), as a suggestion by fellow GitHub users (3), and in recognition of the feature’s utility in conveying valuable information (2). In the following, we analyze the remaining valid responses.
@@
Question Q (, see Figure 5) aimed to assess whether participants consider the presence of personal badges an indicator of their own coding skills in general. Most of the respondents either disagreed () or disagreed strongly (. Then, questions Q-Q () asked participants to report on the effects that badges can have on collaboration with other developers. The results are also reported in Figure 5. Although the distribution of responses is varied, overall the figure shows that about half of the respondents feel that badges cannot be an indicator of others’ ability and that they are not influenced by badges’ presence during collaboration with others. Question Q asked the respondents to justify their previous responses. They were critical of the current implementation of badges. Participant P found that “these badges can be achieved too easily, not like in Stack Overflow,” and P that there are too few of them (“we all have almost the same badges”); others added that they fail to adequately capture developers’ experience – e.g., “I have 30+ years of development experience and almost none is visible through [badges]” (P), “I have lots of experience and almost no badges” (P). Furthermore, P answered with a provocative question: “Are developers with these badges better developers? Vice versa and more critical: Is a developer bad just because they don’t have some of these badges?”
@@
Next, we examine how respondents feel about specific badges. Question Q () asked participants to indicate the most relevant badges. The answers are clustered around two main groups. The first group contains answers (15) from users who feel that none of these badges can tell us anything useful about others. The other group of remaining answers (26) instead believes that the badges give information about who unlocks them. Specifically, 6 respondents mentioned the badge Galaxy Brain (awarded to those who have accepted answers in project discussions) because it reflects “how serious [one is] in the community” (P). A group of 11 participants mentioned the Starstruck badge (unlocked by receiving project stars) because it gives tangible “evidence of recognition and interest by the community” (P). The remaining respondents (9) highlighted the importance of Pair Extraordinaire (awarded to those who coauthor merged pull requests) because the badge may reflect “a positive attitude [toward] collaboration with others” (P).
@@
Finally, question Q () asked participants to indicate any missing badge, in addition to those available. The suggestions converged on the following ideas for new badges that would display: i) the years of work in a project, ii) how many repositories you are a maintainer of, iii) how many repositories you have contributed to, iv) more badges in general, and v) more badges that are challenging to unlock.